OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1).

OpenStudy (anonymous):

find the slope then use point slope formula

OpenStudy (anonymous):

slope: -1 - (-6) / -2 - 3 = 5/-5 = -1 y = -6 -1(x-3) y = -6 -1x +3 y = -x -3